OpenStudy (anonymous):

Kurt's truck uses 10 gallons of gas to travel 120 miles. He has 5 gallons in the tank. How much more gas will he need to drive 276 miles?

OpenStudy (anonymous):

21 more gallons.

OpenStudy (anonymous):

10gals=120mi 120/10=12miles per gallon 12*5=60miles w/5 gallons 276-60=216 miles 216/12=18gallons he needs 18 more gallons of gas.

OpenStudy (anonymous):

thank you @star4120